Convert a speed of 20.4 miles per second into its equivelant in kilometers per hour.
Answer by ankor@dixie-net.com(12701)

Convert a speed of 20.4 miles per second into its equivelant in kilometers per hour.
:
It's useful to know the conversion factor of mph to km/h: Multiply by 1.609
You know there are 3600 seconds in 1 hour
:
Convert miles/sec to mph:
:
Mph = 20.4 * 3600 = 73,440 mph
:
Convert mph to km/h
1.609 * 73440 = 118,165 km/h